OpenStudy (anonymous):

Solve -6(5-2z)+ 5z = 8z - 3(16-3z)

OpenStudy (anonymous):

-30+12z+5z=8z-48+9z -30+17z=17z-48 -30=-48 so this equation produces an inconsistent result and the equation is invalid.